# Fast-dLLM: Training-free Acceleration of Diffusion LLM by Enabling KV Cache and Parallel Decoding

### v1

@@ 3.2 Key-Value Cache for Block-Wise Decoding
Figure 3: Heatmaps of Key-Value Activation Cosine Similarity Across Inference Steps in LLaDA-Instruct. (a) Cosine similarity heatmap for the prompt block, averaged over all prompt tokens. (b) Cosine similarity heatmap for the last block, averaged over all tokens in the last block (used to represent suffix tokens, as the last block always belongs to the suffix before its own decoding). In both (a) and (b), high similarity is observed near the diagonal ($i\approx j$), indicating that Key-Value activations at adjacent inference steps within a block are highly similar. The red boxed regions highlight this effect, supporting the use of an approximate block-wise KV Cache: cached activations from previous steps can be safely reused during block decoding with minimal loss in accuracy. The DualCache strategy, which additionally caches suffix tokens, further demonstrates negligible differences in activations during block decoding, enabling greater acceleration with competitive accuracy
@@ 3.3 Confidence-Aware Parallel Decoding
1:$p_{\boldsymbol{\theta}}$, prompt $p_{0}$, answer length $L$, blocks $K$, block size $B$, steps per block $T$, threshold $\tau$, use_DualCache
@@
9:         Unmask all $i$ in $[s,e)$ with $c^{i}\geq\tau$, always unmask max $c^{i}$ // Parallel decoding
10:         if all $x^{[s,e)}$ unmasked then
11:              break
12:         end if
13:     end for
14:     Update KV cache: if use_DualCache: prefix & suffix; else: prefix. // Cache Update
15:end for
16:return $x$
@@ 4.1 Experimental Setup
We evaluate Fast-dLLM on two recent diffusion-based language models: LLaDA nie2025largelanguagediffusionmodels and Dream dream2025 . Benchmarks include four widely-used datasets—GSM8K, MATH, HumanEval, and MBPP—to assess performance across diverse reasoning and code generation tasks. We also test under varying generation lengths to evaluate scalability and robustness.
@@ 4.2 Main Results: Performance and Speed
Figure 4: Impact of Cache Block Size on Accuracy and Throughput. The orange line illustrates the effect of varying cache block size on throughput, while the blue line depicts the corresponding impact on accuracy.
@@
Importantly, these efficiency gains are achieved with negligible impact on accuracy. Across all benchmarks and settings, the accuracy of our accelerated methods remains within 1–2 points of the backbone, and in several cases, accuracy is even slightly improved. This demonstrates that the speedup comes at almost no cost to task performance, ensuring reliability for practical deployment. We also observe that longer sequences, which are common in few-shot and code generation scenarios, benefit proportionally more from our caching and parallelization techniques due to greater opportunities for cache reuse and batch computation.
Furthermore, the improvements generalize across model architectures (LLaDA and Dream) and task types (math reasoning, program synthesis, etc.), confirming that Fast-dLLM is a practical and broadly applicable framework for accelerating masked diffusion-based language models.
@@ 4.3 Ablations and Analysis
Finally, we evaluate our Confidence-Aware Parallel Decoding method against fixed token-per-step baselines on GSM8K (Figure 5). Our adaptive strategy consistently outperforms fixed baselines across key metrics: it delivers higher accuracy at comparable or reduced number of function evaluations (NFE) and generates more tokens per step on average while closely tracking accuracy. In the rightmost panel, the dynamic method approaches or exceeds the accuracy of the 1-token (non-parallel) baseline, but with much greater throughput. The result demonstrates the effectiveness of Confidence-Aware Parallel Decoding, offering practical advantages.

### v2

@@ 3.2 Key-Value Cache for Block-Wise Decoding
Figure 3: Heatmaps of Key-Value Activation Cosine Similarity Across Inference Steps in LLaDA-Instruct. (a) Cosine similarity heatmap for the prompt block, averaged over all prompt tokens. (b) Cosine similarity heatmap for the last block, averaged over all tokens in the last block (used to represent suffix tokens, as the last block always belongs to the suffix before its own decoding). In both (a) and (b), high similarity is observed near the diagonal ($i\approx j$), indicating that Key-Value activations at adjacent inference steps within a block are highly similar. The red boxed regions highlight this effect, supporting the use of an approximate block-wise KV Cache: cached activations from previous steps can be safely reused during block decoding with minimal loss in accuracy. The DualCache strategy, which additionally caches suffix tokens, further demonstrates negligible differences in activations during block decoding, enabling greater acceleration with competitive accuracy.
@@ 3.3 Confidence-Aware Parallel Decoding
Building on this theorem, we propose a practical factor-based parallel decoding strategy as an extension of the threshold strategy that adaptively selects how many tokens to decode in parallel based on the confidence levels. Concretely, given the model’s marginal confidence estimates for $n$ tokens in a block, we sort these confidences and select the largest $n$ such that $(n+1)(1-c^{(n)})<f$, where $f$ is a fixed decoding factor hyperparameter and $c^{(n)}$ is the $n$-th highest confidence. At each step, the top-$n$ tokens are decoded in parallel. This formulation mirrors the bound in Theorem 1 and ensures that decoding only proceeds when the marginal confidence is sufficiently high to approximate the joint decoding reliably. In contrast to the static threshold-based strategy, factor-based decoding dynamically controls the degree of parallelism in a theoretically grounded manner.
@@
[…]
@@
9:         if strategy == threshold then
10:              Unmask all $i$ in $[s,e)$ with $c^{i}\geq\tau$, always unmask max $c^{i}$
11:         else if strategy == factor then
12:              Sort $c^{i}$ in descending order as $(c^{(1)},c^{(2)},...,c^{(m)})$
13:              Find largest $n$ such that $(n+1)(1-c^{(n)})<f$
14:              Unmask top-$n$ tokens, always unmask the max $c^{i}$
15:         end if
16:         if all $x^{[s,e)}$ unmasked then
17:              break
18:         end if
19:     end for
20:     Update KV cache: if use_DualCache: prefix & suffix; else: prefix. // Cache Update
21:end for
22:return $x$
@@ 4.1 Experimental Setup
Figure 4: Impact of Cache Block Size on Accuracy and Throughput. The orange line illustrates the effect of varying cache block size on throughput, while the blue line depicts accuracy.
We evaluate Fast-dLLM on two recent diffusion-based language models: LLaDA nie2025largelanguagediffusionmodels , LLaDA-1.5 zhu2025llada15variancereducedpreference and Dream dream2025 . Benchmarks include four widely-used datasets—GSM8K, MATH, HumanEval, and MBPP, to assess performance across diverse reasoning and code generation tasks. We also test under varying generation lengths to evaluate scalability and robustness.
In addition, we extend our evaluation to LLaDA-V you2025llada , a multimodal variant of LLaDA tailored for vision-language reasoning tasks. For this, we use two challenging multimodal benchmarks: MathVista and MathVerse, which require solving math problems grounded in complex visual scenes.
@@ 4.2 Main Results: Performance and Speed
Importantly, these efficiency gains are achieved with negligible impact on accuracy. Across all benchmarks and settings, the accuracy of our accelerated methods remains within 1–2 points of the backbone, and in several cases, accuracy is even slightly improved. This demonstrates that the speedup comes at almost no cost to task performance, ensuring reliability for practical deployment. We also observe that longer sequences, which are common in few-shot and code generation scenarios, benefit proportionally more from our caching and parallelization techniques due to greater opportunities for cache reuse and batch computation. We also evaluate an advanced version, LLaDA-1.5, which achieves consistently stronger accuracy and comparable or higher throughput across benchmarks (Table LABEL:tab:llada_1.5).
In addition to text-only models, we evaluate Fast-dLLM on the multimodal LLaDA-V using the MathVista and MathVerse datasets, which require complex vision-language reasoning. As shown in Table 9, LLaDA-V shows a strong sensitivity to block size, with accuracy dropping by over 8% when reducing from 96 to 8 on MathVista. To address this, we retain a full block length and apply refresh-based updates instead of small-block caching. This yields up to $\times9$ speedup with minimal accuracy degradation (Table 3). On MathVerse, accuracy is even slightly improved under Fast-dLLM, demonstrating the broad applicability of our method to multimodal reasoning tasks.
Furthermore, the improvements generalize across model architectures (LLaDA and Dream), task types (math reasoning, program synthesis), and modalities (text and vision), confirming that Fast-dLLM is a practical and broadly applicable framework for accelerating masked diffusion-based language models.
Table 3: Performance and Speedup Comparison of LLaDA-V on MathVista and MathVerse. Each benchmark includes results from Full Steps, Half Steps, and Fast-dLLM. Fast-dLLM significantly improves throughput (highlighted), with minimal accuracy loss.
Metric MathVista MathVerse
Full Steps Half Steps Fast-dLLM Full Steps Half Steps Fast-dLLM
Accuracy (%) 59.2 59.7 56.6 28.5 28.3 28.6
Throughput (Speedup) 2.84 () 5.56 (1.96×) 28.2 (9.9×) 2.75 () 5.17 (1.88×) 23.3 (8.5×)
@@ 4.3 Ablations and Analysis
We evaluate our Confidence-Aware Parallel Decoding method against fixed token-per-step baselines on GSM8K (Figure 5). Our adaptive strategy consistently outperforms fixed baselines across key metrics: it delivers higher accuracy at comparable or reduced number of function evaluations (NFE) and generates more tokens per step on average while closely tracking accuracy. In the rightmost panel, the dynamic method approaches or exceeds the accuracy of the 1-token (non-parallel) baseline, but with much greater throughput. The result demonstrates the effectiveness of Confidence-Aware Parallel Decoding, offering practical advantages.
Factor Decoding vs. Fixed Token-per-Step Strategies
We further compare our factor-based parallel decoding approach with fixed token-per-step baselines on GSM8K (Figure 8) and with the threshold-based strategy (Table 11). Across a range of factor values, our method consistently achieves competitive or higher accuracy with fewer inference steps. As the factor increases, the number of tokens decoded per step grows steadily, reducing iteration count while maintaining performance. Compared to the threshold strategy, factor decoding achieves similar accuracy but significantly higher throughput by adaptively controlling decoding granularity. We also analyze parallel token counts across decoding step at Appendix C.4.
Decoding Efficiency Analysis and Limitations
As discussed in Section C.5, PrefixCache significantly accelerates diffusion-based LLMs like LLaDA with up to 5×5$\times$5 × throughput improvement in compute-bound scenarios compared to LLaDA. At smaller batch sizes, PrefixCache achieves throughput comparable to or even exceeding that of autoregressive models like LLaMA. However, as batch sizes grow, PrefixCache struggles to match LLaMA, which transitions from memory-bound to compute-bound performance. This reflects a general challenge for diffusion-based LLMs, which tend to incur higher computational overhead due to full attention operations during decoding.
@@ B.3 Impact of Dynamic Threshold Settings
B.4 Multimodal Generation with LLAda-V
To qualitatively analyze the effectiveness of our Fast-dLLM framework in multimodal scenarios, we conduct a visual case study where the model is tasked with generating a highly detailed image description. As illustrated in Figure 6, both the baseline model and our Fast-dLLM are given the same visual input and user prompt: “Please describe the image in detail.”
Figure 6: Comparison between the baseline and Fast-dLLM on a visual description task. Fast-dLLM produces a comparable and faithful image caption in a fraction of the decoding time.
The baseline model requires 63.0 seconds to complete the generation, producing a detailed and poetic description of the rural landscape. It highlights elements such as the weathered wooden barn, the soft pink sky, and the tranquil atmosphere.
In contrast, our Fast-dLLM completes the task in just 6.8 seconds—a nearly 10$\times$ speedup—while maintaining rich visual detail. It further enhances the description with additional grounding (e.g., “gray shingles on its roof”, “touch of tranquility”), reflecting a strong alignment with both appearance and mood cues from the image. Notably, the generated caption retains compositional depth and stylistic fluency, illustrating the model’s ability to balance fluency and factuality even under diffusion-based parallel decoding.
This case highlights how LLAda-V with Fast-dLLM decoding enables high-quality vision-language generation at significantly improved efficiency, paving the way for faster and more interactive multimodal applications.
Appendix C Experiment Details
C.1 Further Experiments with LLaDA-V
Table 9: Effect of block length on performance (MathVista, 48 Steps)
Block Length 4 8 16 32 96
Accuracy (%) 51.2 50.7 51.8 52.3 59.7
Throughput (tok./s) 6.1 6.2 5.5 5.5 5.6
Table 10: MathVista Performance with Fast-dLLM at different refresh intervals (block length = 96)
Refresh Interval 2 4 8 16 32
Accuracy (%) 59.2 59.2 58.2 57.1 56.6
Throughput (tok./s) 15.9 19.5 21.1 25.2 28.2
In Table 9, we investigate how the choice of block length affects the performance of LLaDA-V on MathVista under a fixed decoding length of 48 steps. The results show that the model achieves the highest accuracy with a block length of 96. However, when reducing the block size to 8 or 4, the accuracy drops significantly by over 8%.
Given this sensitivity to block length, we choose not to break the output into small blocks for updating caches individually. Instead, we keep the block length fixed at 96 and adopt a refresh-based strategy: the cache is updated only every $r$ decoding steps using the most recent full block. As shown in Table 10, increasing the refresh interval leads to consistent gains in throughput—from 15.9 tokens/s at interval 2 to 28.2 tokens/s at interval 32. While accuracy drops slightly with larger intervals, it remains above 56.6%, suggesting that aggressive refresh scheduling can yield substantial speedups with only minor performance degradation.
C.2 Performance Comparison between Threshold and Factor Strategy
Table 11: Performance comparison between Threshold and Factor confidence-aware decoding on GSM8K and MATH benchmarks with generation lengths of 256 and 512. Each block shows accuracy (top row) and throughput with speedup (bottom row). Factor decoding provides favorable trade-offs in most settings.
Benchmark Gen. Len Threshold Factor
256 78.5 77.5
GSM8K (5-shot) 54.4 (8.1×) 78.5 (11.7x)
512 77.2 74.8
35.3 (11.0×) 47.1 (14.7x)
256 33.2 32.0
MATH (4-shot) 51.7 (5.7×) 78.3 (8.6x)
512 36.0 35.2
47.1 (5.9×) 64.6 (8.1x)
We compare the performance of our threshold-based and factor-based confidence-aware parallel decoding strategies on GSM8K and MATH benchmarks (Table 11). While the threshold strategy achieves marginally better accuracy in most settings (e.g., 78.5% vs. 77.5% on GSM8K with 256 tokens), the factor strategy demonstrates substantially superior throughput performance.
Specifically, factor decoding achieves 1.4-1.5× higher throughput than threshold decoding across all settings. On GSM8K with 256 tokens, factor decoding reaches 78.5 tokens/sec (11.7× speedup) compared to 54.4 tokens/sec (8.1× speedup) for threshold decoding. This throughput advantage becomes even more pronounced on longer generation tasks—for GSM8K with 512 tokens, factor decoding attains 47.1 tokens/sec while threshold only achieves 35.3 tokens/sec.
The results demonstrate that factor decoding offers a compelling trade-off: it sacrifices minimal accuracy (typically 1-3%) in exchange for significant throughput improvements (40-50% higher). This makes factor decoding particularly attractive for latency-sensitive applications where the slight accuracy reduction is acceptable. The consistent pattern across both benchmarks and generation lengths validates the robustness of the factor strategy’s theoretical foundation, which adaptively controls parallelism based on the confidence bound $(n+1)\epsilon<f$.
Figure 7: Average number of tokens generated at each decoding step. Blue line shows the mean token count, and the shaded area denotes the 95% confidence interval.
C.3 Comparison between LLaDA and LLaDA-1.5
We compare the performance of LLaDA and its enhanced version LLaDA-1.5 across both GSM8K (5-shot) and MATH (4-shot) benchmarks under two generation length settings (256 and 512 tokens), as shown in Table LABEL:tab:llada_1.5. Each cell reports accuracy and decoding throughput (in tokens per second), along with the relative speedup over the greedy baseline.
Across GSM8K settings, LLaDA-1.5 consistently improves accuracy over the original LLaDA, achieving a notable +2.2% absolute gain at 256-token generation and +3.2% at 512-token generation. Furthermore, it maintains strong decoding efficiency, with throughput reaching 59.4 tokens/sec at 256 tokens, improving upon LLaDA’s 54.1 tokens/sec under the same setting.
On the MATH benchmark, accuracy between the two versions remains comparable. However, LLaDA-1.5 slightly improves throughput at 256 tokens (53.7 vs. 51.7) while incurring a mild efficiency regression at the 512-token setting (41.1 vs. 47.1). This suggests that while LLaDA-1.5 introduces enhancements beneficial for shorter or moderate decoding contexts, longer sequences may require further optimization.
Overall, LLaDA-1.5 consistently provides either superior accuracy or better decoding speed across settings, demonstrating better performance-efficiency trade-offs and highlighting the benefit of incorporating adaptive improvements on top of the base LLaDA architecture.
Table 12: Performance comparison between LLaDA and LLaDA-1.5. Each cell presents the accuracy and the decoding throughput in tokens per second with relative speedup to the LLaDA baseline (bottom row, blue: tokens per second/orange: relative speedup).
Benchmark Gen Length     LLaDA (Fast-dLLM) LLaDA 1.5 (Fast-dLLM)
GSM8K (5-shot) 256     78.5 80.7
    54.1 (8.1$\times$) 59.4 (8.9$\times$)
512     77.2 80.4
    35.3 (11.0$\times$) 33.0 (10.3$\times$)
MATH (4-shot) 256     33.2 32.6
    51.7 (5.7$\times$) 53.7 (5.9$\times$)
512     36.0 35.1
    47.1 (5.9$\times$) 41.1 (5.1$\times$)
C.4 Analysis of Parallel Token Counts across Decoding Steps
Figure 8: (a) GSM8K (5-shot) accuracy across different factor values using our factor-based decoding strategy. Numbers above each point indicate the average number of tokens decoded per step. The dashed lines show the accuracy of the baseline method with 2 or 4 tokens per step, and the non-parallel (1 token/step) baseline. (b) The corresponding number of inference steps needed under each factor setting. Our method generally requires significantly fewer steps than fixed-step baselines. (c) Accuracy versus average number of tokens decoded per step on GSM8K (5-shot). Our factor-based decoding achieves better accuracy-efficiency trade-offs compared to baselines. The red “Selected” point represents the setting chosen in our main results.
To better understand the behavior of factor-based parallel generation, we analyze the average number of tokens generated at each decoding step. Specifically, we collect statistics from all intermediate steps of the sampling process and compute the average number of tokens generated in parallel per step. The results are visualized in Figure 7, along with a 95% confidence interval indicating cross-sample variability.
As shown in Figure 7, the average number of tokens generated in parallel gradually increases during the early to middle stages of decoding, peaking roughly between step 30 to step 60. After this peak, the parallelism tends to slightly decline toward the end of generation. This suggests that the model becomes more confident in generating outputs during the mid-decoding phase, allowing it to produce more tokens simultaneously. Toward the final steps, the decoding process tends to become more conservative, reducing the number of tokens produced at each step.
The shaded confidence interval reveals greater variance in later decoding steps, indicating instability and inconsistent generation behavior across samples. This is expected since tail-end decoding steps tend to handle only a few remaining tokens required to complete the output, and the number of remaining tokens could differ widely among different samples (e.g., due to early completion or padding).
These observations are important for understanding how decoding efficiency can be optimized: increasing parallelism during high-confidence phases (middle steps) offers computational savings, while conservative behavior near boundaries maintains quality.
C.5 Throughput Comparison under Varying Batch Sizes
Figure 9: Throughput comparison between PrefixCache, LLaDA, and LLaMA under different generation lengths and batch sizes. All models are evaluated on an NVIDIA A100 GPU with the prefill length fixed at 256.
All experiments are conducted on an NVIDIA A100 GPU, with the prefill length fixed to 256 tokens. The generation length is varied among 16, 32, and 64 tokens, and batch sizes range from 1 to 32. This setup reflects realistic deployment scenarios, allowing the evaluation of decoding efficiency under diverse conditions.
It should be noted that parallel decoding allows multiple tokens to be generated simultaneously affected by dummy input tokens. To ensure fairness, we focus solely on the acceleration provided by caching techniques.
PrefixCache is designed as an acceleration mechanism for LLaDA, a diffusion-based LLM, and successfully boosts the throughput significantly. Figure 9 shows that PrefixCache achieves consistent improvements across all batch sizes and generation lengths, making it particularly suited for scenarios with smaller generation lengths and larger batch sizes. For instance, with a generation length of 16 and batch size of 32, PrefixCache achieves a throughput of over 211 tokens/s, significantly outperforming the native LLaDA which reaches only 43 tokens/s, demonstrating nearly 5×5$\times$5 × improvement.
While LLaDA exhibits limited scalability with increasing batch sizes—its throughput plateaus after batch size 8—this limitation is inherent to diffusion-based LLMs, which are compute-bound by nature. In contrast, LLaMA, an autoregressive (AR) model, benefits greatly from large batch sizes. As the batch size increases, LLaMA shifts from being memory-bound to compute-bound, allowing it to achieve high absolute throughput at larger batch settings.
These results highlight the practical advantages of PrefixCache in accelerating compute-bound diffusion models like LLaDA, especially for latency-critical and high-throughput applications. Furthermore, the scalability and efficiency provided by PrefixCache bridge the gap between diffusion-based LLMs and AR models like LLaMA, showcasing its importance for large-scale deployment settings.